Felix drove 800 miles in x hours on
Monday.
a) Write a rational expression for his average speed.
b) On Tuesday he drove for 6 hours at the same average speed. Write a rational expression for his distance on
Tuesday.

Answer by midwood_trail(310)

Felix drove 800 miles in x hours on
Monday.
a) Write a rational expression for his average speed.
b) On Tuesday he drove for 6 hours at the same average speed. Write a rational expression for his distance on
Tuesday.
The distance formula is: distance = rate x time (in hours)
A rational expression is another way of saying fraction.
PART A
D = rt
800 = rx
800/x = r
PART B
D = rt
D = (800/x)(6)
D = 4800/x